INFRARED SPECTRA AND AB INITIO POTENTIAL ENERGY SURFACES FOR ACETYLENE COM-PLEXED TO Mg ATOMS AND CLUSTERS

\underline{DAVID T. MOORE}, ROGER E. MILLER, Department of Chemistry, University of North Carolina, Chapel Hill, NC 27599.

We have recently developed a technique to measure high-resolution infrared spectra of molecules attached to metal atoms and clusters. Using this technique, the spectrum of the CH stretching band of the acetylene-Mg van der Waals complex has been measured in superfluid liquid ^{4}He droplets. A pure b-type band was observed, indicating that the vibrationally averaged structure of the complex is T-shaped, with the Mg atom binding to the π-cloud of the acetylene. Spectral bands corresponding to acetylene-Mg_{n} (n=2-4) complexes were also observed. Structural interpretations based on these spectra will be discussed. High level [CCSD(T)/6-311++G(3df,3pd)] ab initio potential energy surfaces were also constructed for these complexes, and structural predictions based on these calculations will be compared with the experimentally determined structures.